# GRAF: Graph Retrieval Augmented by Facts for Legal Question Answering

### v1

@@ Appendix D Models
The BERT models we used in our work are the RoBERT (masala-etal-2020-robert), a Romanian BERT model trained on general domain and jurBERT (masala2021jurbert) a Romanian BERT model trained on legal domain. After attempting multiple LLM models and comparing their preliminary performances on this task, we proposed the best performing LLMs, of reasonable sizes for our resources for the conducted experiments. We used Flan-T5 (raffel2020exploring), XL 3B and XXL 11B variants, and Mistral 7B (jiang2023mistral). These LLMs are instruction tuned, and we opted for this kind of LLMs for their better performances on instruction following and target tasks (wei2022finetuned), and we could leverage their initial performance on further fine-tuning. For the GAT model, we used 6 attention heads as we noticed experimentally that this value represents an equilibrium in our view between the number of average non-zero entries and computational demands when the GAT is initialized and tested given inputs aiming to potentially ameliorate the dying gradient phenomenon from the null entries. For KG construction and claim extraction we used the Mixtral-8x7B-Instruct LLM (jiang2024mixtral) quantized to 8 bits using the int8 algorithm (dettmers20228bit) implemented in the bitsandbytes library. Even though this model is fairly large, we used it in the KG construction and claim extraction process just as a trustworthy mean which will be more likely to correctly extract the entities and the relations. More lightweight solutions can of-course be built by training a smaller language model for this task or by in fact distilling (hinton2015distilling; gu2023knowledge) Mixtral to a Small Language Model (SLM) for instance. Mixtral did not contribute in any way to helping the rest of our framework in taking the right choice, it had a very clear and definite role in our algorithm which can be easily replaced with any other lightweight solution which we could not perform due to the lack of available data in this sense for the Romanian language.
Appendix E Evaluation Metric
@@
As metric, we used the score that a model would receive on the given test, which is equivalent to the accuracy of the model on the task. No extra points are given if the model miss-predicts correct answers or does not include all correct answers. We used this metric to emphasize the actual test performances of the models. Moreover, we argued that the dataset is balanced and actually cared for comparative analysis and thus we considered this metric will suffice in order to not overload the results sections with more numbers.
Appendix F Detailed Evaluation
@@
Tables 14, 14, 14 show all of our extensive evaluation conducted on different backbone models and different prompts. P1 and P2 refer to the best prompts for Mistral and FLAN-T5 models respectively. Our approach surpasses all of the baseline combinations by using jurBERT-large (masala2021jurbert) as our backbone encoder model. Moreover, it is completely above encoder models in all settings.
@@
Figure 6: The token distribution for the JuRo dataset for Flan-T5 and Mistral models.
@@
Table 11: A detailed view of the JuRo dataset.
@@
Table 12: Results for Promotion exams
@@
Table 13: Results for Entrance exams
@@
Table 14: Results for Bar exams.
Appendix G Romanian Prompts
@@
Appendix H Translated Prompts

### v2

@@ Appendix E Language Models
The BERT models we used in our work are the RoBERT (Masala et al., 2020), a Romanian BERT model trained on general domain and jurBERT (Masala et al., 2021) – a Romanian BERT model trained on the legal domain. After attempting multiple LLM models and comparing their preliminary performances on this task, we found the best-performing LLMs, of reasonable sizes for our resources for the conducted experiments. We used Flan-T5 (Raffel et al., 2020), XL 3B and XXL 11B variants, and Mistral 7B (Jiang et al., 2023). These LLMs are instruction tuned, and we opted for this kind of LLMs for their better performances on instruction following and target tasks (Wei et al., 2022). We could leverage their initial performance on further fine-tuning. For the GAT model, we used 6 attention heads as we noticed experimentally that this value represents an equilibrium in our view between the number of average non-zero entries and computational demands when the GAT is initialized and tested given inputs aiming to potentially ameliorate the dying gradient phenomenon from the null entries. For KG construction and claim extraction, we used the Mixtral-8x7B-Instruct LLM (Jiang et al., 2024) quantized to 8 bits using the int8 algorithm (Dettmers et al., 2022) implemented in the bitsandbytes library. Even though this model is relatively large, we used it in the KG construction and claim extraction process just as a trustworthy mean which will be more likely to correctly extract the entities and the relations. More lightweight solutions can be built by training a smaller language model for this task or by distilling (Hinton, 2015; Gu et al., 2023) Mixtral to a small language model (SLM), for instance. Mixtral did not contribute in any way to helping the rest of our framework in making the right choice. It had a very clear and definite role in our algorithm, which can be easily replaced with any other lightweight solution which we could not perform due to the lack of available data in this sense for the Romanian language.
Appendix F Evaluation Metric
@@
As for the evaluation metric, we used the score that a model would receive on the given test, equivalent to the model’s accuracy on the task. No extra points are given or taken if the model miss-predicts correct answers or does not include all correct answers. We used this metric to emphasize the actual test performances of the models. Moreover, we argued that the dataset is balanced and cared for comparative analysis, and thus, we considered this metric to suffice to not overload the results sections with more numbers.
Appendix G Detailed Evaluation
@@
Tables 15, 15, 15 show all of our extensive evaluations conducted on different backbone models and different prompts. P1 and P2 refer to the best prompts for Mistral and FLAN-T5 models, respectively. Our approach surpasses all baseline combinations by using jurBERT-large (Masala et al., 2021) as our backbone encoder model. Moreover, it is completely above encoder models in all settings.
@@
Figure 6: The token distribution for the JuRO dataset for Flan-T5 and Mistral tokenizers.
@@
Table 12: A detailed view of the JuRO dataset regarding the distribution among legal domains and split into train/test/validation sets.
@@
Table 13: Detailed results for promotion exams.
@@
Table 14: Detailed results for entrance exams.
@@
Table 15: Detailed results for bar exams.
Appendix H Romanian Prompts
@@
[…]
Appendix I Translated Prompts
@@
Claim Graph Extraction - LLM Prompt Extract all entities and all relationships between entities from the legal text based on the example. At the end add STOP. You will answer with triplets of the form: (entity;relation;entity). The triplets are separated on lines. Each triplet relationship will be entered separately. Entities can be institutions, organizations, persons, functions, documents, courts and others. Text: (1) An assets investigation commission, hereinafter referred to as the investigation commission, shall operate in addition to each court of appeal, consisting of: a) 2 judges from the court of appeal, designated by its president, one of whom shall act as president, b) a prosecutor from the prosecutor’s office operating under the court of appeal, designated by the chief prosecutor of this prosecutor’s office. (2) The president and members of the investigation commission shall be designated for a period of 3 years. During the same period and by the same persons, 3 alternates will also be appointed, who will replace the holders in the event that they, for legal reasons, cannot take part in the work of the investigation commission. (3) The investigation commission has a secretary, appointed by the president of the court of appeal from among the clerks of this court. Entity;Relationship;Entity: (court of appeal;shall operated in addition to;assets investigation commission) (assets investigation commission;referred to as;investigation commission) (investigation commission;consisting of;2 judges) (2 judges;designated by;president of the court of appeal) (investigation commission;consisting of;prosecutor) (prosecutor;from;prosecutor’s office operating under the court of appeal) (prosecutor;designated by;chief prosecutor) (president of the investigation commission;designated for a period of;3 years) (members of the investigation commission;designated for a period of;3 years) (3 alternates;appointed by;the president of the court of appeal) (3 alternates;appointed by;the chief prosecutor) (3 alternates;designated for a period of;3 years) (3 alternates;will replace the holders if they cannot take part in the work of the investigation commission on;the heads) (investigation commission;has;a secretary) (a secretary;appointed by among the clerks of;the president of the court of appeal) STOP Text: {text} Entity;Relationship;Entity: